convert to decimal notion.
8.31 x 10^7= _________ (simplify the answer.)

Answer by Edwin McCravy(20054)

                    8.31×107 =

   8.31×10×10×10×10×10×10×10 =

      83.1×10×10×10×10×10×10 =

          831.×10×10×10×10×10 =

            8310.×10×10×10×10 =

              83100.×10×10×10 =
  
                831000.×10×10 =

                  8310000.×10 =
                 
                    83100000. =

                    83100000

Can you figure out a shortcut?  There is a way that takes just 
one step instead of nine.  Examine what happens up there every 
time we multiplied by 10. What happened to the decimal?  How many 
places did it move?  If you can't figure out the shortcut right
away, then do it the above way.  Sooner or later you will figure 
out the decimal-moving way to do it all in one step.

Edwin